OpenStudy (anonymous):

On a roller coaster the first hill has a height of 30m while the second hill has a height of 15m . If the coaster has a speed of 2m/s at the top of the first hill, what is its speed at the top of the second hill if 2% of the energy is lost to friction

OpenStudy (anonymous):

total energy E = K + P ; K is kinetic energy ; P is potential energy K = 1/2 * m * v^2 P = m * g * h So for the first hill, figure out the total energy. Then decrease that total energy by 2%. Then calculate the energy at the second hill. You can get speed from that.